The table shows the performance of a selection of 113 stocks after one year. (Take S to be the set of all stocks represented in the table. If a stock stayed within 20% of its original value, it is classified as "unchanged".)

Companies Total
Pharmaceutical
P
Electronic
E
Internet
I
Increased
V
12 5 16 33
Unchanged
N
33 0 14 47
Decreased
D
12 5 16 33
Total 57 10 46 113

Compute

n(PN').




What does this number represent?

n(PN') is the number of stocks that were either pharmaceutical stocks, or were unchanged after a year (or both).n(PN') is the number of stocks that were either not pharmaceutical stocks, or were not unchanged after a year (or both).     n(PN') is the number of stocks that were either pharmaceutical stocks, or were not unchanged after a year (or both).n(PN') is the number of stocks that were pharmaceutical stocks and were not unchanged after a year.n(PN') is the number of stocks that were not pharmaceutical stocks and were unchanged after a year.
Answer #1

n(P U N') = 113 - P(P N) = 113 - 33 = 80

n(PN') is the number of stocks that were either pharmaceutical stocks, or were not unchanged after a year (or both).